OpenStudy (anonymous):

How many 2-letter "words" consist of two different letters arranged in alphabetical order? (Any two letters together is considered a "word.")

OpenStudy (anonymous):

Should it be arranged in alphabetical order?

OpenStudy (anonymous):

like ab, bc, cd and so on?

OpenStudy (anonymous):

Question is really confusing.

OpenStudy (anonymous):

with a you get 25 with b you get 24 with c you get 23 with y you get 1 with z you get 0 So you get 1+2 +3 + .... +25= 25(26)/2=325